A painting covers 44 square inches one side of the painting is 3 inches more than twice the other side find the dimensions of the painting
Found 2 solutions by checkley77, stanbon:
Answer by checkley77(12569)

L=2W+3
LW=44
(2W+3)W=44
2W^2+3W-44=0
(2W+11)(W-4)=0
W-4=0
W=4 ANS.
L=2*4+3
L=8+3
L=11 ANS.
PROOF:
11*4=44
44=44

Answer by stanbon(48541)

A painting covers 44 square inches one side of the painting is 3 inches more than twice the other side find the dimensions of the painting
-----------------------------
side = "x"
side = 2x+3
--------------------
Equation:
x(2x+3) = 44
2x^2 + 3x - 44 = 0
---
x = [-3 +- sqrt(9 -4*2*-44)]/2
---
x = [-3 +- sqrt(361)]/2
---
Positive solution:
x = [-3 + 19]/2
x = 8 cm (one of the sides)
2x+3 = 19 (the other side)
=================================
Cheers,
Stan H.